Can someone just check this to see if I am correct?? Answer in slope int. form.
A line through the points (-3,4) and (-4,6)
first I found the slope (I THINK)
6-4/-4+3= slope of -2
y-4=-2(x+3)
y-4=-2x-6
y=-2x-6+4
y=-2x-2

Found 2 solutions by funmath, AnlytcPhil:
Answer by funmath(2873)

You're right give yourself a high five!
Happy Calculating!

Answer by AnlytcPhil(1116)

Can someone just check this to see if I am correct?? 
Answer in slope int. form. A line through the points
(-3,4) and (-4,6)
first I found the slope (I THINK)

(6-4)/(-4+3) = slope of -2
y-4=-2(x+3)
y-4=-2x-6
y=-2x-6+4
y=-2x-2

Let's check to see if it's correct.

Substitute (x, y) = (-3,4) and see if that checks.  Then substitute
(x, y) = (-4,6) to see if that also checks. If both check, then you
have to be right:

Substituting (x, y) = (-3,4)

y = -2x - 2
4 = -2(-3) - 2
4 = 6 - 2
4 = 4

So one point checks. Let's check the other point

Substituting (x, y) = (-4,6)

y = -2x - 2
6 = -2(-4) - 2
6 = 8 - 2
6 = 6

That checks too, so you are right.

Edwin